OpenStudy (anonymous):

The sides of a triangular lot are represented by x, 3x, and 3x + 2. Find each side if the perimeter is 632 feet. The lengths of the sides of the triangle are,__,__,and__.

OpenStudy (anonymous):

how did you get 5x+2

OpenStudy (mayankdevnani):

sry

OpenStudy (mayankdevnani):

7x+2=632 x=630/7=90

OpenStudy (mayankdevnani):

1=90,2=270,3=272

OpenStudy (mayankdevnani):

@Claire4christ ok

OpenStudy (mayankdevnani):

@ajprincess right or not

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

thnx...

OpenStudy (ajprincess):

ya it is right:)

OpenStudy (mayankdevnani):

thnx....

OpenStudy (ajprincess):

yw:)